Roberto's mom takes him to the toy store. There are 6 types of green balls, 8 types of blue balls and 5 types of red balls. If Roberto's mom allows him to get one of each color, how many options does Roberto have?
Answer by ewatrrr(24785)

 
Hi,
6 types of green balls, 8 types of blue balls and 5 types of red balls.
If Roberto's mom allows him to get one of each color, how many options does Roberto have?
6C1*8C1*5C1 = 6*8*5 = 240 options